OpenStudy (abhisar):

Can you factorise \(\sf \large \frac{4AB}{(A+B)(A+B)}\) to \(\large\sf\frac{4B}{A}\)

OpenStudy (anonymous):

what?

OpenStudy (abhisar):

one min !

OpenStudy (anonymous):

factorization or factoring is the decomposition of an object into a product of other objects

OpenStudy (abhisar):

\(\sf \large \frac{4AB}{(A+B)(A+B)}\) to \(\large\sf\frac{4B}{A}\)

OpenStudy (anonymous):

if its for explanding brackets \[a(b + c) \], the expanded version is \[ab + ac\]

OpenStudy (vishweshshrimali5):

Uhm It is not correct: \[\large{\cfrac{4AB}{(A+B)^2} \ne \cfrac{4B}{A}}\]

OpenStudy (abhisar):

yep..that's what i was thinkin !

OpenStudy (anonymous):

for that denominator must equal a^2 ,, , so no

OpenStudy (abhisar):

may be i should post the original question. But it's physics

OpenStudy (abhisar):

Find the ratio x:y if x= \(\sf \large \frac{CAB}{d^2}\) y= \(\sf \large \frac{C\times (A+B)/2\times (A+B)/2}{d^2}\)

OpenStudy (vishweshshrimali5):

\[\large{x:y = \cfrac{4AB}{(A+B)^2}}\]

OpenStudy (vishweshshrimali5):

Doesn't appear to be a Physics question @Abhisar O.o :)

OpenStudy (anonymous):